MATH 2090  Problem Set #4 (On Ch.14)

York University

Faculties of Pure and Applied Science, Arts, Atkinson MATH 2090. Problem Set #4. Posted March 9, 2002

Due in the Course Box by 2pm. March 20, 2002

Section N

1. Do from text p.300 and onwards: 14.10, 14.14, 14.20, 14.26. Also do

2.

ST ⊢ (∃x)(⟨x, y⟩ = z) ∧ (∃x)(⟨x, w⟩ = z) ⇒ y = w

Page 1  G. Tourlakis